# TokMem: Tokenized Procedural Memory for Large Language Models

### v1

@@ Appendix A Decoupled Embedding for TokMem
Figure 5: Overview of Decoupled TokMem embeddings, which learns separate memory matrices for address of memories and generation steering.
@@
In the standard TokMem formulation, each memory token embedding ${\bm{m}}_{i}\in\mathbb{R}^{d}$ is shared across two roles: (1) addressing for memory routing and (2) steering for generation. We consider a decoupled (DC) variant that separates these functions into two embedding matrices:
$M^{\text{addr}}\;=\;\big[{\bm{u}}_{1},\ldots,{\bm{u}}_{l}\big]\in\mathbb{R}^{l\times d},\qquad M^{\text{steer}}\;=\;\big[{\bm{s}}_{1},\ldots,{\bm{s}}_{l}\big]\in\mathbb{R}^{l\times d}.$ (8)
@@
Here, $M^{\text{addr}}$ provides address embeddings at the output layer. When a memory token is predicted, the model produces a distribution over indices $i$ according to $M^{\text{addr}}$. The chosen index $i$ is then used to retrieve the corresponding steering embedding ${\bm{s}}_{i}$ from $M^{\text{steer}}$, which is injected into the input sequence and influences subsequent generation.
@@
$\mathcal{L}\big({\bm{a}};M^{\text{addr}},M^{\text{steer}}\big)\;=\;-\sum_{i>k}\log\Pr\big(a_{i}\;\big|\;{\bm{a}}_{<i};M^{\text{addr}},M^{\text{steer}}\big),$ (9)
where $k$ denotes the query length. During optimization, only $M^{\text{addr}}$ and $M^{\text{steer}}$ are updated; the backbone remains frozen. In addition, the renormalization treatment introduced in Section 2.2 is only applied to the address embeddings in $M^{\text{addr}}$.
This decoupled formulation provides a clean separation of functionality: routing is handled via $M^{\text{addr}}$, while steering is controlled by $M^{\text{steer}}$. While conceptually clean, our experiments do not show consistent improvements over the coupled formulation, particularly on larger models, where embedding capacity is sufficient to jointly support both roles.
Appendix B Details for Compositional Memory Recall
B.1 Details of Adaptation Phase
In compositional scenarios, the model should not only recall individual procedures but also compose them to solve multi-step queries. To prepare TokMem for such use, we construct a held-out auxiliary training set of $50$ tools (5,0005{,}000 samples) following Section 3.3. The backbone is then fine-tuned for one epoch on this set using LoRA, jointly with the temporary memory embeddings, before the adapted weights are merged and frozen.
The intuition for this adaptation phase to let the LLM learn to align its routing and generation behavior with compositional memory recall. After adaptation, the temporary embeddings are discarded, while the adapted backbone is retained for inference with new tasks. This procedure provides a general inductive bias for modular composition, enabling it to generalize to new tools and procedures without further retraining.
Algorithm 1 summarizes this lightweight procedure. Temporary memory embeddings are inserted into the input sequence, the loss is optimized jointly over memory and response tokens, and once the backbone has adapted, the temporary memory bank is discarded.
@@ B.1 Details of Adaptation Phase
1:Pretrained backbone $f_{\theta_{0}}$, adaptation traces $\mathcal{D}_{\text{adapt}}$ from held-out procedures
2:Initialize backbone $\theta\leftarrow\theta_{0}$ and temporary memory embeddings $\mathcal{M}$
@@
5:  Insert $\mathcal{M}$ into sequence; forward pass with $f_{\theta}$
6:  Compute loss $\mathcal{L}$ on memory and response tokens
@@
10:Discard temporary memory $\mathcal{M}$ and freeze backbone $\theta$
11:return adapted backbone $f_{\theta}$
B.2 Analysis of Unfreezing LLM Backbone for TokMem
Figure 6: Comparison between Freezing and unfreezing the backbone. Allowing the backbone to update when adding new tool memories causes severe forgetting. Freezing preserves prior tools while enabling new ones.
We further examine the importance of freezing the backbone when adding new tool memories, reflecting real-world usage where procedural knowledge grows incrementally over time. This setting contrasts with recent approaches (Wang et al., 2025) that compress tool usage into virtual tokens by post-training the backbone. While such methods improve retrieval efficiency at scale, they rely on modifying backbone parameters, which hinders continual adaptability and risks overwriting prior knowledge.
As shown in Figure 6, unfreezing the backbone during TokMem adaptation leads to severe forgetting of previously learned tools, consistent with catastrophic interference in continual learning. By contrast, freezing the backbone preserves prior capabilities while allowing new tool memories to be incorporated without loss, highlighting TokMem’s advantage for incremental adaptation. Notably, unfreezing offers no meaningful performance gains after the initial training round, suggesting that TokMem strikes an effective balance between performance and continual adaptation.
Appendix C Additional Analysis on Memory Placement
Table 6: Comparison of prefix tuning vs. TokMem condition embedding on GSM8K with two different size of Llama models. TokMem achieves higher compliance with required output formats and stronger exact-match accuracy than prefix tuning, especially in low-data regimes.
@@ Appendix C Additional Analysis on Memory Placement
We have stress-tested the effect of memory token placement (prefix vs. infix) with randomly generated queries and with varying length of memory tokens in Section 3.4. We now turn to the GSM8K math reasoning dataset (Cobbe et al., 2021), to evaluate generalization and training efficiency.
@@
Our experiments run on Llama 3.2 1B and 3B as backbone models and compare prefix tuning against TokMem under two training setups: using only $20\%$ of the training set that represents a low-data regime, or the full dataset. We report two evaluation metrics.
@@
  • Compliance measures whether the model follows the required answer format, i.e., producing the final answer after the delimiter “####”. This metric isolates the recall of procedural memory from the reasoning abilities already present in the backbone models.
@@
  • Exact Match (EM) measures the correctness of the final answer after standard normalization (e.g., removing commas or extraneous symbols).
@@
As shown in Table 6, TokMem significantly outperforms prefix tuning, particularly in the low-data setting. With only $20\%$ of the data, prefix-tuning fails to provide meaningful results, yielding zero compliance and EM on the 1B model and underperforming on the 3B model. By contrast, TokMem achieves near-perfect compliance and substantially higher EM scores across both backbones. When trained on the full dataset, prefix tuning improves considerably, yet TokMem continues to deliver stronger compliance and higher EM, underscoring its superior data efficiency and more reliable procedural control.
Appendix D Details of Datasets
D.1 Details of Super-Natural Instruction
We sample 1,0001{,}000 English tasks from the SNI dataset, where each task is labeled with a task ID and a short descriptive name. The full list of sampled tasks is provided in Table 7. Tasks are introduced to the model sequentially in ascending order of their IDs (e.g., the model first sees task 1, then task 2, and so on).
After training on the first $k$ tasks, we save a checkpoint and evaluate performance on the test sets of all $k$ tasks encountered so far. This simulates a continual learning setup where the model is expected to acquire new procedures while retaining previously learned ones. Once the model has been trained on all 1,0001{,}000 tasks, it should be able to perform all of them without forgetting earlier tasks.
D.2 Details of Function Calling dataset
For evaluating compositional memory recall, we sample $50$ tools from the APIGen dataset (Liu et al., 2024b). The list of tools and their corresponding descriptions is provided in Table 7.
For each tool, we collect $50$ query–call pairs, some of which may involve multiple calls to the same tool. This yields a total of $50\times 50=2{,}500$ samples representing the non-compositional use of tools. To avoid data leakage, we split these samples into training and test sets with a 9:19{:}1 ratio.
On top of this, we synthesize complex queries by combining calls across different tools. These multi-step queries require the model to invoke multiple tools in sequence. We cap the number of synthesized samples at 5,0005{,}000 for training and $500$ for testing.
@@ D.2 Details of Function Calling dataset
Table 7: Details of the sampled tools from the APIGen dataset, including their names and descriptions.

### v2

@@ A.2 Details of Decoupled Embedding for TokMem
Figure 5: Overview of the decoupled variant of TokMem embeddings, which learn separate memory matrices for memory indexing and generation steering.
@@
In the standard TokMem formulation, each memory token embedding ${\bm{m}}_{i}\in\mathbb{R}^{d}$ is shared across two roles: (1) indexing for memory routing and (2) condition to steer for generation. We consider a decoupled (DC) variant that separates these functions into two embedding matrices:
$M^{\text{index}}=({\bm{u}}_{1};\dots;{\bm{u}}_{l})\in\mathbb{R}^{l\times d},\qquad M^{\text{steer}}=({\bm{s}}_{1};\dots;{\bm{s}}_{l})\in\mathbb{R}^{l\times d}.$ (7)
@@
As seen in Figure 5, $M^{\text{index}}$ provides index embeddings at the output layer. The model produces a distribution over indices $i$ according to $M^{\text{index}}$ when a memory token is predicted. The chosen index $i$ is then used to retrieve the corresponding steering embedding ${\bm{s}}_{i}$ from $M^{\text{steer}}$, which is then appended into the sequence and conditions subsequent generation.
@@
$\mathcal{L}({\bm{a}};M^{\text{index}},M^{\text{steer}})=-\sum_{i>k}\log\Pr(a_{i}\mid{\bm{a}}_{<i};M^{\text{index}},M^{\text{steer}}).$ (8)
where $k$ denotes the query length. During optimization, only $M^{\text{index}}$ and $M^{\text{steer}}$ are updated, while the backbone remains frozen. In addition, the renormalization treatment introduced in Section 2.2 is applied only to the index embeddings $M^{\text{index}}$.
This decoupled formulation provides a clean separation of functionality: routing is handled via $M^{\text{index}}$, while steering is controlled by $M^{\text{steer}}$. However, our experiments do not show consistent improvements over the coupled formulation, particularly on larger models where their larger embedding capacity is sufficient to jointly support both roles. We therefore advocate for the simple yet effective TokMem (without DC).
A.3 Details of Function-Calling Dataset
To evaluate compositional memory recall, we sample $50$ tools from the APIGen dataset (Liu et al., 2024b). The list of tools and their corresponding descriptions is provided in Table 7.
For each tool, we collect $50$ query–call pairs, some of which may involve multiple calls to the same tool. We categorize the use of the same tool as non-compositional tool use, yielding a total of $50\times 50=2{,}500$ samples.
Additionally, we augment the dataset by synthesizing complex queries by combining calls across different tools. These multi-step queries require the model to invoke multiple tools in sequence. To avoid data leakage, we split these samples into training and test sets, capping the synthesized samples at 5,0005{,}000 for training and $500$ for testing.
A.4 Details of Adaptation Phase
In compositional scenarios, the model should not only recall individual procedures but also compose them to solve multi-step queries. To prepare TokMem for such use, we construct a held-out auxiliary training set comprising an additional $50$ tools (5,0005{,}000 samples) that is disjoint from the tool set detailed in Appendix A.3. The backbone is then fine-tuned for one epoch on this set using LoRA, jointly with the temporary memory embeddings.
The intuition behind this adaptation phase is to let the base LLM learn to use the embeddings with compositional memory recall, rather than memorizing on the tool descriptions themselves. After adaptation, the temporary embeddings are discarded, while the adapted backbone is retained for training on new tools. This procedure provides a general inductive bias for modular composition, enabling the model to generalize to new tools and procedures without further retraining.
Algorithm 1 summarizes this lightweight procedure. Temporary memory embeddings are inserted into the input sequence, the loss is optimized jointly over memory and response tokens, and the temporary memory bank is discarded once the backbone has adapted.
@@ A.4 Details of Adaptation Phase
1:Pre-trained backbone $f_{\theta_{0}}$, adaptation traces $\mathcal{D}_{\text{adapt}}$ from held-out procedures
2:Initialize the backbone $\theta\leftarrow\theta_{0}$ and temporary memory embeddings $\mathcal{M}$
@@
5:  Insert $\mathcal{M}$ into the sequence; run a forward pass with $f_{\theta}$
6:  Compute the loss $\mathcal{L}$ on memory and response tokens
@@
10:Discard the temporary memory $\mathcal{M}$ and freeze the backbone $\theta$
11:return the adapted backbone $f_{\theta}$
Appendix B Additional Results
B.1 Results on Compositional Generalization
Table 5: TokMem generalizes to longer tool-call chains at test time, significantly outperforming fine-tuning in zero-shot multi-step settings (trained with 1 call and tested with 2–4 calls).
Test Time
Train Maximum Calls Method 2 calls 3 calls 4 calls Avg.
1-call Fine-tuning 34.9 21.3 14.1 23.4
TokMem 60.3 54.3 48.9 54.5 (+31.1)
2-call Fine-tuning 86.2 78.8 64.8 76.6
TokMem 82.0 81.8 82.3 82.0 (+5.4)
3-call Fine-tuning 86.9 85.5 80.3 84.2
TokMem 86.8 84.0 84.7 85.2 (+1.0)
4-call Fine-tuning 87.9 86.6 82.9 85.8
TokMem 85.9 86.7 88.3 86.3 (+0.5)
We observe that TokMem provides clear advantages in compositional generalization over fine-tuning. Table 5 reports Argument F1 when the Llama 3B model is evaluated on queries that require more function calls than those observed during training.
Notably, when trained solely on single-call data, TokMem achieves much stronger performance than fine-tuning when evaluated on test data with 2244 calls. This demonstrates that memory tokens trained for atomic procedures can be effectively composed at test time, enabling strong zero-shot generalization to multi-step behavior.
As the training regime is expanded to include more calls (e.g., up to $3$ or $4$), the performance gap narrows, but TokMem remains competitive or slightly ahead across configurations. These results suggest that TokMem naturally supports compositionality, enabling flexible chaining of learned procedures without requiring task-specific fine-tuning.
B.2 Analysis of Norm Inflation for Newer Memories
Figure 6: $\ell_{2}$ norm of newly added memory tokens. Each round introduces 10 new tool memories. Error bars indicate the standard deviation across the 10 tokens added in each round.
We analyze the $\ell_{2}$ norm of the learned memory embeddings when tools are introduced sequentially with Llama 3.2 3B. We follow the setup in Figure 4 by training TokMem for $5$ rounds and adding $10$ tools per round without our renormalization treatment detailed in Section 2.4. Figure 6 shows that newly added memory tokens gradually increase their $\ell_{2}$ norms, which leads to competition with existing frozen tokens in the softmax used for memory routing.
B.3 Analysis of Unfreezing LLM Backbone for TokMem
Figure 7: Comparison between freezing and unfreezing the backbone. Allowing the backbone to update when adding new tool memories causes severe forgetting, whereas freezing preserves prior tools while enabling new ones.
We further examine the importance of freezing the backbone when adding new tool memories, reflecting real-world usage where procedural knowledge grows incrementally over time. This setting contrasts with recent approaches (Wang et al., 2025) that aim to compress tool use description by post-training the backbone. Such methods rely on modifying backbone parameters, which hinders continual adaptation and risks overwriting prior knowledge.
As shown in Figure 7, We confirm that unfreezing the backbone during TokMem adaptation leads to severe forgetting of previously learned tools. By contrast, our choice of freezing the backbone meaningfully preserves prior capabilities while allowing new tool memories to be incorporated without loss, highlighting TokMem’s advantage for incremental adaptation. This analysis suggests that TokMem strikes an effective balance between performance and continual adaptation.
B.4 Additional Analysis on Memory Placement
Table 6: Comparison of prefix tuning and TokMem conditioning embeddings on GSM8K with two different Llama model sizes. TokMem achieves higher compliance with required output formats and stronger exact-match accuracy than prefix tuning, especially in low-data regimes.
@@ B.4 Additional Analysis on Memory Placement
In Section 3.4, we have stress-tested the memorization with different memory token placement (prefix vs. infix) by varying numbers of memory tokens. We now turn to the GSM8K math reasoning dataset (Cobbe et al., 2021) to evaluate generalization and training sample efficiency.
@@
Our experiments use Llama 3.2 1B and 3B as backbone models and compare prefix tuning against TokMem under two training setups: using only $20\%$ of the training set (a low-data regime) or the full dataset. We report two evaluation metrics.
@@
  • Compliance measures whether the model follows the required answer format, i.e., producing the final answer after the delimiter “####”. This metric isolates procedural-memory recall from the reasoning abilities already present in the backbone.
@@
  • Exact Match (EM) measures the correctness of the final answer after standard text normalization (e.g., removing commas or extraneous symbols).
@@
As shown in Table 6, TokMem significantly outperforms prefix tuning, particularly in the low-data setting. With only $20\%$ of the data, prefix tuning fails to provide meaningful results, yielding zero compliance and EM on the 1B model and underperforming on the 3B model. By contrast, TokMem achieves near-perfect compliance and substantially higher EM scores across both backbones. When trained on the full dataset, prefix tuning improves considerably, yet TokMem continues to deliver stronger compliance and higher EM, showing its superior data efficiency and more reliable procedural control.
Appendix C Future Work
Our experiments are conducted on the research-oriented SNI and APIGen datasets, which provide a controlled environment for analyzing atomic and compositional recall. While these settings demonstrate the feasibility and effectiveness of one-token procedural memory without backbone training, they do not fully capture the diversity and ambiguity of real-world procedures.
In particular, richer forms of composition, such as interleaving function calls from APIGen with NLP tasks from SNI, are bottlenecked by TokMem’s reliance on manually curated, procedurally decomposed datasets. To bridge this gap, a future direction is the automation of query–procedural decomposition. By leveraging highly capable teacher LLMs to synthesize interleaved training trajectories, we can automatically map complex queries to discrete procedural tokens, providing a scalable alternative to manual annotation.
Beyond automated data synthesis, additional future directions include incorporating reinforcement learning to improve generalization over complex compositional structures, and enabling personalization by allowing users to attach custom, user-specific memory banks while keeping the backbone frozen. These extensions pave the way for scalable, compact, and user-adaptive memory systems in large language models.
Appendix D Use of Large Language Models
We used ChatGPT as a general-purpose assistant to improve the writing of our paper, including grammar, readability, and clarity. Additionally, we used it to search for potentially missing related work, which we then manually read and is discussed in the paper. All ideas, analyses, and conclusions presented in this paper are our own, and we take full responsibility for the content.
@@ Appendix D Use of Large Language Models
Table 7: Details (names and descriptions) of the sampled tools from the APIGen dataset.